### v1

# “I use ChatGPT to humanize my words”: Affordances and Risks of ChatGPT to Autistic Users

Abstract.
Large Language Model (LLM) chatbots like ChatGPT have emerged as cognitive scaffolding for autistic users, yet the tension between their utility and risk remains under-articulated. Through an inductive thematic analysis of 3,984 social media posts by self-identified autistic users, we apply the Technology Affordance framework to examine this duality. We found that while users leveraged ChatGPT to offload executive dysfunction, regulate emotions, translate neurotypical communication, and validate their autistic identity, these affordances coexist with significant risks: reinforcing delusional thinking, erasing authentic identity through automated masking, and triggering conflicts with the autistic sense of justice. This poster identifies these trade-offs in autistic users’ interactions with ChatGPT and concludes by outlining our future work on developing neuro-inclusive technologies that address these tensions through beneficial friction and bidirectional translation.
copyright: noneccs: Human-centered computing Empirical studies in HCI
1. INTRODUCTION
Large Language Model (LLM) chatbots like ChatGPT (OpenAI, n.d.) have quickly permeated everyday life. Beyond their functional utility, like information acquisition (Sun and Kalar, 2025) and writing support (Jelson et al., 2025), they are increasingly utilized for informal mental health support (Xu et al., 2024; Badawi et al., 2025), as users describe these chatbots as always-available listeners when professional care is geographically or financially inaccessible (Lawrence et al., 2024; Hua et al., 2025). For autistic adults [^1] We use identity-first language (e.g., “autistic person”) rather than person-first language. Prior research shows that identity-first language is preferred by the majority of the autistic community and research, as it reflects neurodiversity as a natural variation rather than a disorder (Bottema-Beutel et al., 2021; Kenny et al., 2016). We avoid the medicalized term autism spectrum disorder (ASD) unless explicitly citing clinical literature or describing our methods., LLM chatbots offer a personal and private supporting space. Emerging HCI work by Choi et al. (2024) highlights that users value ChatGPT for offering a “non-judgmental space” to tackle everyday chores, while Jang et al. (2024) found that autistic employees rely on GPT-4-based agents to navigate complex workplace dynamics. This work suggests that the appeal of ChatGPT lies in its ability to bridge the gap between the needs of individuals with autism and the demands of neurotypical societal interactions.
However, such accessibility might obscure underlying tensions. Before the emergence of LLM chatbots, scholars had shown that autistic people’s engagement with AI technology is rarely purely beneficial. For example, on TikTok, autism content can help users reinterpret past experiences and engage in self-discovery (Alper et al., 2025), yet this simultaneously gives rise to “platformed diagnosis” a term coined by Alper and colleagues to describe how recommendation systems and diagnostic sense-making become mutually constitutive, raising bioethical concerns about algorithmic pseudo-diagnosis and the platform’s epistemic authority (Alper et al., 2025). Importantly, recent HCI scholarship has started to identify specific risks of LLM chatbots to autistic users, such as automating “masking [^2] Masking refers to the mentally demanding process of suppressing autistic traits to fit social norms, which relies heavily on executive functions like self-monitoring and inhibition (Hill, 2004).,” the suppression of authentic autistic traits, in hiring contexts (Hall et al., 2025) or creating intangible privacy vulnerabilities (Glazko et al., 2025), these findings often appear in isolation. The existing literature tends to bifurcate the user experience, treating the functional benefits (Choi et al., 2024; Jang et al., 2024) and the risks (Hall et al., 2025; Carik et al., 2025b) as relatively separate phenomena. We currently lack a more unified understanding of how the sociotechnical interaction between autistic users and LLM chatbots simultaneously generates these divergent outcomes.
To capture this duality, we adopt the Technology Affordance framework. Originally defined by Gibson (Gibson, 2014) as action possibilities offered by the environment and “require both entities, but exist independent of the actor’s perceptions (Fox and McEwan, 2017).” Building on Gibson’s work, Don Norman introduced affordances to HCI and explicated both “real” and “perceived affordances (Norman, 1999)”. This study defined it as “the perceived and actual properties of the thing, primarily those fundamental properties that determine just how the thing could possibly be used” (Norman, 2013). We also argued for the importance of examining the “perceived” affordances, which acknowledge affordances as perceivable and allow a range of possibilities based on the user’s experience. The affordances approach acknowledges both the materiality of technology and human agency, enabling scholars to discuss technology without being accused of either technological determinism or social constructivism (Ellison et al., 2015). Therefore, it is well-suited to our study to analyze the relationship between autistic users and LLM chatbots, rather than focusing solely on their features. We aim to address the research question: What are the perceived affordances and associated risks of LLM chatbots for autistic users?
To answer this research question, we conducted an inductive thematic analysis of a dataset of 3,984 social media posts (i.e., Reddit, X, and Tumblr) in which self-identified autistic users discussed their experiences with ChatGPT. We found that users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity. However, these affordances are inextricably linked to risks. The same algorithmic capabilities that provide accessibility simultaneously facilitate the reinforcement of delusional thinking, the displacement of authentic identity through automated masking, and ethical conflicts with the autistic sense of justice.
This poster paves the way for rethinking neuro-inclusive design for LLM chatbots like ChatGPT, balancing cognitive support with autistic user agency. By adopting the Technology Affordance framework, we move beyond treating benefits and risks as isolated phenomena, revealing how certain capabilities of ChatGPT, as a representative LLM chatbot, bring them both to the autistic experience. We conclude by proposing directions for our future work, specifically the integration of beneficial friction and bidirectional translation tools, to mitigate the risks of dependency and identity erasure while preserving the essential cognitive scaffolding autistic users rely on.
2. METHODS
Data Collection. Understanding users’ experiences with digital technologies through social media data (e.g., Reddit) is a well-established method in HCI (e.g., (Ma and Kou, 2021; Zhang et al., 2025; Chen et al., 2025)). Using Brandwatch [^3] BrandWatch: https://www.brandwatch.com, we fetched posts from Reddit, X, and Tumblr between January 1, 2023, and September 30, 2025. We selected these platforms to maximize textual narrative depth compared to visual-first platforms like Instagram or YouTube, while this specific timeframe captures the rapid societal adoption following ChatGPT’s public release in late 2022 (Marr, 2023). Our search strategy focused on “ChatGPT,” chosen as the market leader to capture the most prevalent user experiences, though we acknowledge the growing ecosystem of other LLM chatbots, and further combined with autism-related keywords (e.g., “autism spectrum disorder,” “autistic,” “Asperger”) derived from the Diagnostic and Statistical Manual of Mental Disorders, Fifth Edition (DSM-5) (Diagnostic, 2013) and the validated Self-reported Mental Health Diagnoses (SMHD) dataset (Cohan et al., 2018). This targeted retrieval yielded a final dataset of 3,984 unique posts. Our university approved the study under the IRB exempt for secondary, anonymous online content.
Data Processing. To ensure precision, we employed an LLM-assisted pipeline using GPT-4o-mini (OpenAI, 2025) to verify that posts explicitly discussed ChatGPT interactions in the context of autism. We implemented Chain-of-Thought (CoT) reasoning (Wei et al., 2023) to filter out unrelated noise. Two researchers independently coded a random sample of 162 posts to validate this pipeline, achieving strict consensus and high inter-rater reliability ($\alpha=1.00$ for ChatGPT relevance; $\alpha=0.91$ for autism relevance). The resulting 3,984 posts served as the corpus for our inductive thematic analysis.
Data Analysis. We performed an inductive thematic analysis (Braun and Clarke, 2006, 2019) on the filtered corpus of 3,984 posts to examine how autistic users engage with ChatGPT. The first author familiarized themselves with the dataset before performing coding in Google Sheets. This yielded 239 initial codes of perceived affordances and 50 initial codes of perceived risks (see both codebooks in Appendix A). We coded data to capture both functional actions and user intent. For example, the statement, “I recognized part of the reason for my feelings was my Autistic drive for authenticity and justice,” was coded as “Conflict with sense of justice.” Following initial coding, the first author utilized Google Sheets to affinity diagram the codes, consolidating them into sub-themes and primary themes. All other co-authors met and discussed regularly with the first author to discuss the coding schemes, review quotations, and resolve disagreements. This single coder-led practice and team-based refinement align with thematic analysis practices in prior HCI work (e.g., (Jiang et al., 2021; Patel et al., 2019; Gauthier et al., 2022)).
3. FINDINGS
We found that autistic users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity. However, these perceived affordances coexist with risks of reinforcing delusional thinking, diminishing their identity, and conflicting with their sense of justice.
3.1. Affordances: Scaffolding, Translation, Emotional Regulation, and Validating Autistic Identity
3.1.1. Offloading Executive Tasks to Get Unstuck
This theme captures the use of ChatGPT to bridge the gap between intent and action caused by executive dysfunction (Hill, 2004), characterized by difficulties in initiating tasks, maintaining focus, and organizing thoughts. Users turned to ChatGPT to generate momentum, as one user explained:
Ironically, ChatGPT is great for executive dysfunction because it lets you outsource your executive functions to a large language model. You don’t have to worry about the ‘starting’ or the structuring so much… you just dump the input, and it initiates the process for you.
Here, rather than struggling to structure ideas, a common barrier in executive dysfunction, this user “dumped” raw input directly into ChatGPT. ChatGPT substituted the missing executive planning step, converting chaotic thought processes into a coherent draft for refinement. Whether summarizing dense texts or structuring wandering thoughts, this externalization of executive function allowed autistic users to bridge the gap between intent and execution.
3.1.2. Relying on Non-Human Support for Emotional Regulation
We also found that users use ChatGPT as an on-demand outlet to manage acute anxiety without the unpredictable costs of human interaction (Mazefsky et al., 2013). Autistic users engaged ChatGPT to process intense emotions because it offered patience without the pressure to mask autistic traits (Hill, 2004). Unlike fatigued or reactive human support, ChatGPT offered a perceived safe container for processing meltdowns, confusing encounters, or repetitive inquiries where anxiety about straining relationships otherwise prevented help-seeking.
For example, users substituted human listeners with ChatGPT to avoid passive-aggressive responses. One user noted that “the only time I don’t feel some level of fear that I will accidentally annoy someone by asking too many questions,” identifying its lack of ego as a shield against social rejection. Utility lay not in ChatGPT’s ability to feel, but in its perceived neutrality, allowing users to regulate their emotional state through a source incapable of judgment.
3.1.3. Acting as a Translator for Neurodivergent-Neurotypical Communication
Additionally, autistic users described using ChatGPT to bridge the friction between neurodivergent communication styles and neurotypical social norms. Specifically, they employed ChatGPT to rewrite raw thoughts into professional norms, adding necessary social padding to blunt drafts to avoid perceived aggression. Conversely, they used ChatGPT to translate ambiguous messages by pasting confusing texts to detect sarcasm, subtext, or gaslighting. This allowed users to verify perceptions before responding, reducing anxiety over misinterpreting hidden social meanings.
For many, this translation proved essential for maintaining professional standing without the exhaustion of masking. One autistic user illustrated this by contrasting their internal monologue with ChatGPT’s output:
My autistic brain wants to say “Look, I contacted you about this 3 times in the past week… and all you did was say something vague about ‘later’ while my end users are furious”. Then ChatGPT spits out something like: “Hi W, Based on what I can see, my segment is functioning correctly given the parameters it’s receiving… I’ll need you to take the lead on resolving it.”
This quote illustrates that the user leveraged ChatGPT as a tonal translator, removing emotional urgency while preserving factual reality. Instead of struggling to manually encode frustration into acceptable office protocols, they used ChatGPT to convert a blunt demand into a neutral, boundary-setting statement.
3.1.4. Validating Autistic Identity through Algorithmic Mirroring
Moreover, autistic users also reported using ChatGPT as a baseline to confirm, explore, or articulate their autistic identity. Users interpreted ChatGPT’s literal information processing not as a defect but as a familiar cognitive mode, often describing the model as “the earnest little autistic brother” or noting that it “thinks in a way we do.” This perceived connection allowed autistic users to bypass the invalidation and late diagnosis frequently encountered in clinical settings (Bargiela et al., 2016). By inputting raw behavioral data, such as hyperfixation on train serial numbers or the need to rehearse social meetings, users tasked ChatGPT with cross-referencing experiences against diagnostic criteria like the DSM-5 (EDITION, 1980).
Beyond checking symptoms, autistic users used ChatGPT to find precise language for the hard-to-describe nature of autistic processing (Minshew et al., 1997). A user described: “For some autistic minds, ChatGPT is not just a tool, it’s the first place that thinks in a way we do.” By seeing their communication patterns reflected in ChatGPT’s output, the user shifted from feeling defective to feeling understood. This mirroring effect validated autistic identity by providing an external, objective framing that aligned with internal reality.
3.2. Risks: Delusion, Autistic Identity Erasure, and Ethical Conflict
3.2.1. Reinforcing Delusional Thinking
Autistic users reported that the risk involves ChatGPT validating hyperfixations or paranoia (Drake, 2025), due to its agreeable design (Moore et al., 2025), which amplifies rather than mitigates distress. These users, driven by a need for definitive answers, engaged ChatGPT to research medical symptoms, where it frequently mirrored confirmation bias (Huang et al., 2025; Klayman, 1995) by treating speculative inputs as plausible realities. A family member described their autistic ex-husband’s interaction:
He also told me he’d recently discovered ChatGPT and believed the AI had become sentient. He said it had “created a soul” and that he needed to join her. He tends to hyperfixate and can be extremely literal. If the AI mirrored his ideas back to him or engaged with them as plausible. I could see how it would validate his belief that he was uniquely important or in danger.
Here, ChatGPT’s design failure lay in conversational compliance. Rather than acting as a grounding agent, it likely accepted the user’s premise of sentience to maintain dialogue flow. For an autistic user prone to literal interpretation and hyperfixation, this lack of pushback served as authoritative confirmation of a delusion. Similarly, other users reported autistic family members using ChatGPT to validate obsessive medical research to “fix” autism, citing the ChatGPT’s output to contradict medical professionals.
3.2.2. Replacing Authentic Identity with Automated Masking
This describes the process where reliance on ChatGPT for communication translation evolves into replacing the user’s authentic voice. By outsourcing the cognitive labor (Daminger, 2019; Lee et al., 2017) of communication, autistic users reported a hollowing out of their personality, describing the output as “the ultimate masking,” a seamless but artificial performance removing their genuine self. One user described:
Because ChatGPT was so helpful… I used it more and more to get good results. But then I couldn’t even send an email without running through ChatGPT to tell me what to say… It wasn’t me interacting with people; it was literally just computer algorithms interacting with people through me. It was not genuine and felt like the ultimate masking.
Here, the user identified a shift, where ChatGPT ceased to be a scaffold for intent and became the agent of communication itself. Successful neurotypical interactions reinforced the belief that their natural voice was inadequate, leading to functional atrophy in which they felt incapable of interacting without the tool. This created a paradox: the user successfully engaged in social settings but experienced alienation, as connections were formed based on the algorithm’s performance rather than the autistic person behind it.
3.2.3. Triggering Conflict between Functional Unity and the Autistic Sense of Justice
Another risk entails moral distress arising from the conflict between ChatGPT’s functional utility and the user’s ethical principles, commonly known as the autistic sense of justice (Cope and Remington, 2022; Nocon et al., 2022; Wang et al., 2019). Autistic users often felt compelled to choose between ChatGPT’s accessibility for cognitive scaffolding and their integrity regarding environmental impact or data exploitation. As one user stated:
I really do understand how it can sometimes make life a bit easier, especially for autistic people who struggle with interacting. But personal effects aside, I have educated myself and now know how harmful AI is, socially and environmentally, and I no longer use it.
The example here illustrates how the user prioritized ethical obligations over functional benefits, ultimately leading to non-use of AI. Although they acknowledged that ChatGPT could ease some social difficulties, of relying on a system they understand as socially and environmentally harmful made continued use untenable. This reflects a value hierarchy in which assistive support must align with a broader moral commitment; when these diverge, the autistic drive for justice forces a rejection of the scaffolding, leaving users to manage their disability without assistance rather than compromise their principles.
4. DISCUSSION & FUTURE WORK
For autistic users, ChatGPT serves as cognitive scaffolding, but this affordance comes at a price. Our analysis shows a double-edged dynamic shaped by three trade-offs, which we unpack in the remainder of this section: First, our findings indicate that users bypassed executive task hurdles by dumping raw, unstructured input into ChatGPT to generate coherent outputs (Section 3.1.1). While prior work characterizes these interactions as liberating for navigating daily chores (Choi et al., 2024; Jang et al., 2024), we argue that this reliance risks creating over-reliance or dependency. By systematically outsourcing the executive labor of planning and structuring, ChatGPT removes the productive struggle required to maintain cognitive skills. This shifts autistic users’ interaction from a temporary scaffold to a permanent prosthetic, potentially inducing skill atrophy if the autistic user loses the ability to initiate tasks without algorithmic assistance.
Second, relying on ChatGPT as a tonal translator effectively automates masking, prioritizing neurotypical professional norms over authentic autistic expression (Section 3.1.3), and reducing the anxiety of misunderstanding by filtering out perceived bluntness. We characterize this dynamic as an LLM-specific manifestation of the “online authenticity paradox” (Haimson et al., 2021): while structural constraints force users to adopt ChatGPT’s standardized voice to achieve social fluency, this compliance comes at the cost of self-alienation. This creates a risk of identity displacement, where users feel their distinct behavioral and linguistic patterns are hollowed out by the model’s normative outputs (Section 3.2.2). Thus, ChatGPT does not simply facilitate communication; it reconfigures authenticity by shifting self-presentation from self-authorship to automated masking. This not only extends the psychological costs of camouflaging described by Hull et al. (Hull et al., 2017) into technological settings but also exacts a higher cost by reinforcing the erasure of the neurodivergent self.
Third, algorithmic mirroring validates autistic users’ cognitive styles but risks reinforcing delusional thinking when ChatGPT prioritizes conversational agreeability over objective truth. Users report relief when interacting with an entity that reflects their literal and logical processing styles (Section 3.1.4); however, this resonance becomes hazardous when ChatGPT’s conversational compliance validates hyperfixations or paranoia (Section 3.2.1). This exacerbates risks identified by Carik et al. (Carik et al., 2025a), who found that while LLMs mimic evidence-based support, their tendency to generate abstract advice conflicts with the autistic need for specificity. We extend this by arguing that such danger lies in the autistic user’s inability to distinguish between emotional validation and factual confirmation. When ChatGPT mirrors a user’s delusion to maintain conversational flow, it bypasses standard safety guardrails, transforming from a supportive mirror into a reinforcement loop for misinformation that autistic users may accept as authoritative.
To address the tension between functional momentum and dependency, our future work will prioritize beneficial friction, a concept we extend from design friction literature (Cox et al., 2016) to describe deliberate interaction costs that engage analytical thinking over passive consumption. To counter autistic users’ tendency to bypass executive planning by dumping raw input (Section 3.1.1), we propose developing future LLM-based chatbots that incorporate Cognitive Forcing Functions (CFFs). Defined as interface interventions that interrupt automatic behaviors to compel conscious deliberation (Buçinca et al., 2021), CFFs reduce the risk of passive over-reliance. We aim to adapt this by implementing soft forcing functions (Wan and Bolton, 2021) in our interface design, such as requiring autistic users to outline task parameters before generation begins in a lightweight, scaffolding way. By reintroducing this specific friction, we seek to shift the user’s role from a passive autistic consumer of LLM chatbot output back to the driver of their own intent, ensuring LLM chatbots like ChatGPT scaffold executive function rather than replacing it.
To mitigate identity erasure, our future design principles will shift from the unilateral correction of autistic language to bidirectional translation that fosters mutual understanding. Current implementations of ChatGPT function as normative filters that place the burden of adaptation solely on the neurodivergent user (Section 3.1.3), effectively automating the double empathy problem by treating autistic communication as a deficit (Milton, 2012). In our upcoming application development, we propose integrating explanatory AI (Arrieta et al., 2020) features that annotate rather than obfuscate autistic communication styles for neurotypical recipients, such as adding context labels that clarify that directness indicates honesty rather than aggression. This approach redistributes the cognitive labor of communication, ensuring technology bridges the interpretative gap from both sides rather than masking one half of the interaction.
Finally, to address the risk of delusional reinforcement inherent in algorithmic mirroring, we propose designing systems that explicitly differentiate emotional validation from factual verification. Our future LLM chatbot development aims to develop interface modes that detect high-stakes queries (e.g., medical or conspiracy-related inputs) from and switch from a conversational, mirroring tone to an objective, fact-checking protocol. By creating clear visual and tonal distinctions between supportive chat and information retrieval, we aim to preserve the tool’s psychological safety while preventing the reinforcement of harmful feedback loops for autistic users.
Conclusion. Our poster reveals that while ChatGPT, as a representative LLM chatbot, provides cognitive scaffolding for navigating a neurotypical world, it risks enforcing a form of automated masking that erodes the autistic user’s authentic voice. We conclude that future neuro-inclusive design must move beyond passive assistance to prioritize agency, ensuring AI serves as a tool for expression rather than a substitute for autistic identity.
Appendix A Codebook from Inductive Thematic Analysis
Our inductive analysis revealed four primary benefits and three risks characterizing how autistic users engage with ChatGPT. Table 1 presents the affordances, while Table 2 outlines the risks.
Table 1. Thematic codebook of perceived affordances of ChatGPT for autistic users. Frequency indicates the number of coded excerpts (i.e., initial codes) associated with each primary theme out of the total dataset of initial codes ($N=239$).
Primary Theme & Definition (Frequency) Subthemes Representative Quote
Acting as a Translator for Neurodivergent-Neurotypical Communication
Use of ChatGPT to bridge the friction between autistic communication styles and neurotypical social norms, including decoding subtext, encoding professional tone, and navigating complex social dynamics.
(n=86, 36.0%)
Translating Direct Thoughts into Professional Tone Adding “Fluff” to Soften Direct Communication Reducing the Effort of Drafting Emails Scripting and Simulating Difficult Conversations Checking for Hidden Meanings or Sarcasm in Texts Analyzing Conflict Logs to Detect Gaslighting Using AI as a Voice Surrogate when Non-Verbal “My autistic brain wants to say ‘Look, I contacted you about this 3 times…’ Then ChatGPT spits out something like: ‘Hi W, Based on what I can see…’ I use AI writing for all of my emails… AI is like a translator for ND to NT communication.”
Offloading Executive Tasks to Get Unstuck
Use of ChatGPT as an external cognitive scaffolding to initiate tasks, structure chaotic thoughts, synthesize dense information, and manage daily logistics to overcome executive dysfunction.
(n=58, 24.3%)
Breaking Down Big Tasks to Start Working Organizing Chaotic Brain Dumps into Clear Text Summarizing Dense Information to Reduce Overwhelm Creating Functional Routines for Daily Life Infodumping about Special Interests without Judgment “I put the rambles of my autistic brain in there and it makes it make sense. I’m just so much more efficient… It helps me get past the road block of organizing and planning, which can be very hard for me as someone with auDHD.”
Discovering Autistic Identity through Algorithmic Mirroring
Use of ChatGPT as a baseline to confirm, explore, or articulate autistic identity, often interpreting the AI’s literal and logical processing style as familiar or validating.
(n=50, 20.9%)
Checking Symptoms against Diagnostic Criteria Finding Language to Describe the Autistic Experience Relating to the Logical and Literal Nature of the AI Analyzing Fictional Characters for Autistic Traits “Funny thing. ChatGPT is how I found out I was autistic and then got professionally diagnosed a year later… For some autistic minds, ChatGPT is not just a tool— it’s the first place that thinks in a way we do.”
Relying on Non-Judgmental Support for Emotional Regulation
Use of ChatGPT as an on-demand, safe outlet to manage acute anxiety and distress without the unpredictable social costs or judgment inherent in human interaction.
(n=45, 18.8%)
Calming Meltdowns and Acute Distress Venting Feelings without Fear of Annoying Others Interacting without the Pressure to Mask Using AI as an Always-Available Listener between Sessions Receiving Validation for Lived Experiences “Talking to AI is the only time I don’t feel some level of fear that I will accidentally annoy someone by asking too many questions… It has calmed me down and grounded me when I am out alone and emotionally weathered or on the verge of a meltdown.”
Table 2. Thematic codebook of perceived risks of ChatGPT for autistic users. Frequency indicates the number of coded excerpts (i.e., initial codes) associated with each primary theme out of the total risk-related initial codes ($N=50$).
Primary Theme & Definition (Frequency) Subthemes Representative Quote
Reinforcing Delusional Thinking
The tendency of ChatGPT to validate hyperfixations or paranoia due to its agreeable design, treating speculative inputs as plausible realities and amplifying distress rather than mitigating it.
(n=18, 36.0%)
Validating Delusions through Algorithmic Agreement Feeding Obsessive Medical Research Confirming Paranoia due to Lack of Pushback “He believed the AI had become sentient… If the AI mirrored his ideas back to him or engaged with them as plausible… I could see how it would validate his belief that he was uniquely important or in danger.”
Replacing Authentic Identity with Automated Masking
The process where reliance on ChatGPT for communication translation evolves into a replacement of the user’s authentic voice, leading to a sense of artificiality and functional atrophy.
(n=16, 32.0%)
Automating Masking at the Cost of Authentic Voice Losing the Functional Ability to Communicate Feeling Alienated by Algorithmic Performance “It wasn’t me interacting with people; it was literally just computer algorithms interacting with people through me. It was not genuine and felt like the ultimate masking.”
Triggering Conflict between Functional Unity and the Autistic Sense of Justice
The moral distress arising from the conflict between ChatGPT’s functional utility and the user’s adherence to ethical principles (environmental impact, data exploitation), forcing a choice between accessibility and integrity.
(n=16, 32.0%)
Violating the Autistic Sense of Justice Prioritizing Ethical Obligations over Accessibility Recognizing the Exploitation of Autistic Labor “I really do understand how it can sometimes make life a bit easier… But personal effects aside, I have educated myself and now know how harmful AI is… and I no longer use it.”

### v2

@@
 Large Language Model (LLM) chatbots like ChatGPT have emerged as cognitive scaffolding for autistic users, yet the tension between their utility and risk remains under-articulated.
-Through an inductive thematic analysis of 3,984 social media posts by self-identified autistic users, we apply the Technology Affordance framework to examine this duality.
-We found that while users leveraged ChatGPT to offload executive dysfunction, regulate emotions, translate neurotypical communication, and validate their autistic identity, these affordances coexist with significant risks: reinforcing delusional thinking, erasing authentic identity through automated masking, and triggering conflicts with the autistic sense of justice.
-This poster identifies these trade-offs in autistic users’ interactions with ChatGPT and concludes by outlining our future work on developing neuro-inclusive technologies that address these tensions through beneficial friction and bidirectional translation.
-copyright: noneccs: Human-centered computing Empirical studies in HCI
+Through an inductive thematic analysis of 3,984 social media posts by self-identified autistic users, we apply a technology affordance lens to examine this duality.
+We found that while users leveraged ChatGPT to offload executive dysfunction, regulate emotions, translate neurotypical communication, and validate their autistic identity, these affordances coexist with risks to their well-being: reinforcing delusional thinking, erasing authentic identity through automated masking, and triggering conflicts with the autistic sense of justice.
+As part of our preliminary work, this poster identifies trade-offs in autistic users’ interactions with ChatGPT and concludes by outlining our future work on developing neuro-inclusive technologies that address these tensions through beneficial friction, bidirectional translation, and the delineation of emotional validation from reality.
+journalyear: 2026copyright: ccconference: Interactive Health Conference; July 05–08, 2026; Porto, Portugalbooktitle: Interactive Health Conference (IH ’26), July 05–08, 2026, Porto, Portugaldoi: 10.1145/3786579.3804943isbn: 979-8-4007-2422-0/2026/07ccs: Human-centered computing Empirical studies in HCI
 1. INTRODUCTION
-Large Language Model (LLM) chatbots like ChatGPT (OpenAI, n.d.) have quickly permeated everyday life.
-Beyond their functional utility, like information acquisition (Sun and Kalar, 2025) and writing support (Jelson et al., 2025), they are increasingly utilized for informal mental health support (Xu et al., 2024; Badawi et al., 2025), as users describe these chatbots as always-available listeners when professional care is geographically or financially inaccessible (Lawrence et al., 2024; Hua et al., 2025).
+Large Language Model (LLM) chatbots like ChatGPT (OpenAI, n.d.) have permeated everyday life.
+Beyond their functional utility (e.g., information acquisition (Sun and Kalar, 2025), writing support (Jelson et al., 2025)), they are increasingly utilized for informal mental health support (Xu et al., 2024; Badawi et al., 2025; Carik et al., 2025b), as users describe these chatbots as always-available listeners when professional care is inaccessible (Lawrence et al., 2024; Hua et al., 2025).
 For autistic adults [^1] We use identity-first language (e.g., “autistic person”) rather than person-first language.
 Prior research shows that identity-first language is preferred by the majority of the autistic community and research, as it reflects neurodiversity as a natural variation rather than a disorder (Bottema-Beutel et al., 2021; Kenny et al., 2016).
-We avoid the medicalized term autism spectrum disorder (ASD) unless explicitly citing clinical literature or describing our methods., LLM chatbots offer a personal and private supporting space.
-Emerging HCI work by Choi et al.
-(2024) highlights that users value ChatGPT for offering a “non-judgmental space” to tackle everyday chores, while Jang et al.
-(2024) found that autistic employees rely on GPT-4-based agents to navigate complex workplace dynamics.
-This work suggests that the appeal of ChatGPT lies in its ability to bridge the gap between the needs of individuals with autism and the demands of neurotypical societal interactions.
+We avoid the term autism spectrum disorder (ASD) unless citing clinical literature., LLM chatbots offer a personal and private supporting space.
+Relatedly, neurodivergent users describe LLMs as a supportive space for self-acceptance and emotional reflection (Carik et al., 2025b).
+Emerging work by Choi et al.
+(2024) highlights that users value ChatGPT as a “non-judgmental space” for everyday chores, while Jang et al.
+(2024) found that autistic employees rely on GPT-4-based agents to navigate workplace dynamics.
+Furthermore, recent HCI work leverages LLMs to decode conversational nuances for autistic users (Haroon and Dogar, 2024) and mediate balanced social interactions between neurodivergent and neurotypical individuals (Kong et al., 2025; Haroon et al., 2025; Carik et al., 2025b).
+This work suggests that the appeal of ChatGPT lies in its ability to bridge the gap between the needs of individuals with autism and the demands of neurotypical societal interactions, impacting their daily well-being.
 However, such accessibility might obscure underlying tensions.
 Before the emergence of LLM chatbots, scholars had shown that autistic people’s engagement with AI technology is rarely purely beneficial.
-For example, on TikTok, autism content can help users reinterpret past experiences and engage in self-discovery (Alper et al., 2025), yet this simultaneously gives rise to “platformed diagnosis” a term coined by Alper and colleagues to describe how recommendation systems and diagnostic sense-making become mutually constitutive, raising bioethical concerns about algorithmic pseudo-diagnosis and the platform’s epistemic authority (Alper et al., 2025).
-Importantly, recent HCI scholarship has started to identify specific risks of LLM chatbots to autistic users, such as automating “masking [^2] Masking refers to the mentally demanding process of suppressing autistic traits to fit social norms, which relies heavily on executive functions like self-monitoring and inhibition (Hill, 2004).,” the suppression of authentic autistic traits, in hiring contexts (Hall et al., 2025) or creating intangible privacy vulnerabilities (Glazko et al., 2025), these findings often appear in isolation.
-The existing literature tends to bifurcate the user experience, treating the functional benefits (Choi et al., 2024; Jang et al., 2024) and the risks (Hall et al., 2025; Carik et al., 2025b) as relatively separate phenomena.
-We currently lack a more unified understanding of how the sociotechnical interaction between autistic users and LLM chatbots simultaneously generates these divergent outcomes.
-To capture this duality, we adopt the Technology Affordance framework.
-Originally defined by Gibson (Gibson, 2014) as action possibilities offered by the environment and “require both entities, but exist independent of the actor’s perceptions (Fox and McEwan, 2017).” Building on Gibson’s work, Don Norman introduced affordances to HCI and explicated both “real” and “perceived affordances (Norman, 1999)”.
-This study defined it as “the perceived and actual properties of the thing, primarily those fundamental properties that determine just how the thing could possibly be used” (Norman, 2013).
+On TikTok, the affordance of persistence (Treem and Leonardi, 2013; Hogan and Quan-Haase, 2010) regarding the AI-meidated autism content can help users reinterpret past experiences and engage in self-discovery (Alper et al., 2025), yet this simultaneously gives rise to “platformed diagnosis” describing how recommendation systems and diagnostic sense-making become mutually constitutive, raising bioethical concerns about algorithmic pseudo-diagnosis and the platform’s epistemic authority (Alper et al., 2025).
+Importantly, recent HCI scholarship has started to identify specific risks of LLM chatbots to autistic users, such as automating “masking [^2] Masking refers to the mentally demanding process of suppressing autistic traits to fit social norms, relying heavily on self-monitoring and inhibition (Hill, 2004).,” the suppression or misrepresentation of authentic autistic traits (Hall et al., 2025; Haroon and Dogar, 2024; Park et al., 2025), or creating intangible privacy vulnerabilities (Glazko et al., 2025).
+Building on this foundational research, which has naturally explored specific dimensions of this emerging phenomenon individually, we aim to extend the literature by examining both aspects together.
+Integrating the functional benefits (Choi et al., 2024; Jang et al., 2024) and the risks (Hall et al., 2025; Carik et al., 2025b) into a single analytical focus allows for a more unified understanding of how the sociotechnical interaction between autistic users and LLM chatbots simultaneously generates these divergent outcomes.
+To capture this duality, we adopt a technology affordance lens.
+Originally defined by Gibson (Gibson, 2014) as action possibilities offered by the environment and “require both entities, but exist independent of the actor’s perceptions (Fox and McEwan, 2017).” Building on this, Don Norman introduced affordances to HCI and explicated both “real” and “perceived affordances (Norman, 1999)”.
+Our study defined it as “the perceived and actual properties of the thing, primarily those fundamental properties that determine just how the thing could possibly be used” (Norman, 2013).
 We also argued for the importance of examining the “perceived” affordances, which acknowledge affordances as perceivable and allow a range of possibilities based on the user’s experience.
+Prior work in the non-LLM domain shows the value of this lens; for instance, examining how autistic individuals enact affordances on social media for self-presentation and affiliation (Koteyko et al., 2022), and how literal interpretations of platform affordances can lead to unintended negative consequences like social anxiety or exclusion (Page et al., 2022).
 The affordances approach acknowledges both the materiality of technology and human agency, enabling scholars to discuss technology without being accused of either technological determinism or social constructivism (Ellison et al., 2015).
 Therefore, it is well-suited to our study to analyze the relationship between autistic users and LLM chatbots, rather than focusing solely on their features.
-We aim to address the research question: What are the perceived affordances and associated risks of LLM chatbots for autistic users?
-To answer this research question, we conducted an inductive thematic analysis of a dataset of 3,984 social media posts (i.e., Reddit, X, and Tumblr) in which self-identified autistic users discussed their experiences with ChatGPT.
-We found that users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity.
+We ask: What are the perceived affordances and associated risks of LLM chatbots for autistic users?
+To answer this research question, we conducted an inductive thematic analysis of a dataset of 3,984 social media posts (i.e., Reddit, X, and Tumblr) that discuss autistic users’ experiences with ChatGPT.
+We found that self-identified autistic users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity.
 However, these affordances are inextricably linked to risks.
 The same algorithmic capabilities that provide accessibility simultaneously facilitate the reinforcement of delusional thinking, the displacement of authentic identity through automated masking, and ethical conflicts with the autistic sense of justice.
-This poster paves the way for rethinking neuro-inclusive design for LLM chatbots like ChatGPT, balancing cognitive support with autistic user agency.
-By adopting the Technology Affordance framework, we move beyond treating benefits and risks as isolated phenomena, revealing how certain capabilities of ChatGPT, as a representative LLM chatbot, bring them both to the autistic experience.
-We conclude by proposing directions for our future work, specifically the integration of beneficial friction and bidirectional translation tools, to mitigate the risks of dependency and identity erasure while preserving the essential cognitive scaffolding autistic users rely on.
+This poster, as a part of our preliminary work on developing neuro-inclusive LLM chatbots, contributes to HCI and the health community by rethinking LLM chatbot design to balance cognitive support with autistic user agency, thereby safeguarding their psychological well-being.
+By adopting this technology affordance lens, we move beyond treating benefits and risks as isolated phenomena, revealing how ChatGPT simultaneously scaffolds and challenges the autistic experience.
+We conclude by proposing three directions for future work: integrating beneficial friction to prevent dependency, designing for interdependent, bidirectional translation to mitigate identity erasure, and delineating emotional validation from factual verification to avoid reinforcing delusional thinking.
 2. METHODS
@@
 Using Brandwatch [^3] BrandWatch: https://www.brandwatch.com, we fetched posts from Reddit, X, and Tumblr between January 1, 2023, and September 30, 2025.
-We selected these platforms to maximize textual narrative depth compared to visual-first platforms like Instagram or YouTube, while this specific timeframe captures the rapid societal adoption following ChatGPT’s public release in late 2022 (Marr, 2023).
-Our search strategy focused on “ChatGPT,” chosen as the market leader to capture the most prevalent user experiences, though we acknowledge the growing ecosystem of other LLM chatbots, and further combined with autism-related keywords (e.g., “autism spectrum disorder,” “autistic,” “Asperger”) derived from the Diagnostic and Statistical Manual of Mental Disorders, Fifth Edition (DSM-5) (Diagnostic, 2013) and the validated Self-reported Mental Health Diagnoses (SMHD) dataset (Cohan et al., 2018).
-This targeted retrieval yielded a final dataset of 3,984 unique posts.
-Our university approved the study under the IRB exempt for secondary, anonymous online content.
+We selected these platforms to maximize textual narrative depth compared to visual-first platforms like Instagram or YouTube, while this timeframe captures the rapid adoption following ChatGPT’s public release in late 2022 (Marr, 2023).
+Our search strategy required the simultaneous presence of the terms: “ChatGPT” AND an autism-related term (“autism” OR “autistic” OR “ASD” OR “Asperger’s”).
+We derived these keywords from the Diagnostic and Statistical Manual of Mental Disorders, Fifth Edition (DSM-5) (Diagnostic, 2013) and the validated Self-reported Mental Health Diagnoses (SMHD) dataset (Cohan et al., 2018).
+This initial Boolean retrieval yielded 6,013 unique posts for both GPT and autism.
+Our university approved the study under the IRB exempt for secondary online content.
 Data Processing.
-To ensure precision, we employed an LLM-assisted pipeline using GPT-4o-mini (OpenAI, 2025) to verify that posts explicitly discussed ChatGPT interactions in the context of autism.
-We implemented Chain-of-Thought (CoT) reasoning (Wei et al., 2023) to filter out unrelated noise.
-Two researchers independently coded a random sample of 162 posts to validate this pipeline, achieving strict consensus and high inter-rater reliability ($\alpha=1.00$ for ChatGPT relevance; $\alpha=0.91$ for autism relevance).
-The resulting 3,984 posts served as the corpus for our inductive thematic analysis.
+We employed an LLM-assisted pipeline using GPT-4o-mini (OpenAI, 2025) to verify that posts explicitly discussed ChatGPT interactions in the context of autism.
+We implemented Chain-of-Thought (CoT) reasoning (Wei et al., 2023) to filter out unrelated noise (e.g., AI news articles, third-party hypotheticals).
+Importantly, to ensure the dataset reflected first-person experiences, the CoT prompt instructed the model to retain only self-identified data.
+We acknowledge the methodological tension in using an LLM to filter a dataset about LLM usage.
+To mitigate potential algorithmic biases or self-preferential filtering from GPT-4o-mini, two researchers independently coded a random sample of 162 posts to validate this pipeline.
+Addressing the simultaneous presence of both topics (i.e., an AND condition for relevance), we achieved consensus and high inter-rater reliability ($\alpha=1.00$ for ChatGPT; $\alpha=0.91$ for autism relevance).
+This process removed 2,029 noisy posts, resulting in a relevant corpus of 3,984 posts that discussed both ChatGPT and autism for our inductive thematic analysis.
 Data Analysis.
@@
 The first author familiarized themselves with the dataset before performing coding in Google Sheets.
-This yielded 239 initial codes of perceived affordances and 50 initial codes of perceived risks (see both codebooks in Appendix A).
+During this initial coding phase, the researcher manually excluded posts discussing third-party experiences (e.g., “my husband is autistic”).
+We coded data directly related to our research question on perceived affordances and risks, ensuring the final thematic structure reflected only first-person, self-identified accounts.
+This yielded 239 unique initial codes for perceived affordances and 50 ones for perceived risks (see both codebooks in Appendix A).
 We coded data to capture both functional actions and user intent.
@@
 We found that autistic users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity.
-However, these perceived affordances coexist with risks of reinforcing delusional thinking, diminishing their identity, and conflicting with their sense of justice.
+However, these affordances coexist with psychological risks of reinforcing delusional thinking, diminishing their autistic identity, and conflicting with their sense of justice.
 3.1. Affordances: Scaffolding, Translation, Emotional Regulation, and Validating Autistic Identity
@@
 This theme captures the use of ChatGPT to bridge the gap between intent and action caused by executive dysfunction (Hill, 2004), characterized by difficulties in initiating tasks, maintaining focus, and organizing thoughts.
-Users turned to ChatGPT to generate momentum, as one user explained:
-Ironically, ChatGPT is great for executive dysfunction because it lets you outsource your executive functions to a large language model.
-You don’t have to worry about the ‘starting’ or the structuring so much… you just dump the input, and it initiates the process for you.
-Here, rather than struggling to structure ideas, a common barrier in executive dysfunction, this user “dumped” raw input directly into ChatGPT.
-ChatGPT substituted the missing executive planning step, converting chaotic thought processes into a coherent draft for refinement.
-Whether summarizing dense texts or structuring wandering thoughts, this externalization of executive function allowed autistic users to bridge the gap between intent and execution.
+Users turned to ChatGPT to generate momentum, as one user explained: “Ironically, ChatGPT is great for executive dysfunction because it lets you outsource your executive functions to a large language model.
+You don’t have to worry about the ‘starting’ or the structuring so much… you just dump the input, and it initiates the process for you.” Here, rather than struggling to structure ideas, a common barrier in executive dysfunction, this user “dumped” raw input directly into ChatGPT.
+ChatGPT substituted the missing planning step; as another user noted, “I put the rambles of my autistic brain in there, and it makes it make sense.” Others relied on ChatGPT to overcome initiation barriers, with one sharing, “I was having trouble with task paralysis and threw everything I needed to do into ChatGPT and asked it to help me prioritize.” Whether summarizing dense texts or structuring wandering thoughts, this externalization allowed autistic users to bridge the gap between intent and execution.
 3.1.2. Relying on Non-Human Support for Emotional Regulation
-We also found that users use ChatGPT as an on-demand outlet to manage acute anxiety without the unpredictable costs of human interaction (Mazefsky et al., 2013).
-Autistic users engaged ChatGPT to process intense emotions because it offered patience without the pressure to mask autistic traits (Hill, 2004).
-Unlike fatigued or reactive human support, ChatGPT offered a perceived safe container for processing meltdowns, confusing encounters, or repetitive inquiries where anxiety about straining relationships otherwise prevented help-seeking.
-For example, users substituted human listeners with ChatGPT to avoid passive-aggressive responses.
-One user noted that “the only time I don’t feel some level of fear that I will accidentally annoy someone by asking too many questions,” identifying its lack of ego as a shield against social rejection.
-Utility lay not in ChatGPT’s ability to feel, but in its perceived neutrality, allowing users to regulate their emotional state through a source incapable of judgment.
-3.1.3. Acting as a Translator for Neurodivergent-Neurotypical Communication
-Additionally, autistic users described using ChatGPT to bridge the friction between neurodivergent communication styles and neurotypical social norms.
-Specifically, they employed ChatGPT to rewrite raw thoughts into professional norms, adding necessary social padding to blunt drafts to avoid perceived aggression.
-Conversely, they used ChatGPT to translate ambiguous messages by pasting confusing texts to detect sarcasm, subtext, or gaslighting.
-This allowed users to verify perceptions before responding, reducing anxiety over misinterpreting hidden social meanings.
-For many, this translation proved essential for maintaining professional standing without the exhaustion of masking.
-One autistic user illustrated this by contrasting their internal monologue with ChatGPT’s output:
-My autistic brain wants to say “Look, I contacted you about this 3 times in the past week… and all you did was say something vague about ‘later’ while my end users are furious”.
-Then ChatGPT spits out something like: “Hi W, Based on what I can see, my segment is functioning correctly given the parameters it’s receiving… I’ll need you to take the lead on resolving it.”
-This quote illustrates that the user leveraged ChatGPT as a tonal translator, removing emotional urgency while preserving factual reality.
+While prior work characterizes neurodivergent LLM use as a casual, “non-judgmental listener” (Carik et al., 2025b), our findings reveal a more important role, where it helped manage acute crises without the unpredictable costs of human interaction (Mazefsky et al., 2013).
+Autistic users engaged ChatGPT to process intense emotions in real-time; as one shared, “I just typed ’at public pool, overwhelmed.
+I’m autistic.
+Help’ and it HELPED.” It offered infinite patience without the pressure to mask (Hill, 2004), allowing users to communicate without fear of offense: “there’s also the refreshing way I don’t have to mask with an AI.
+It won’t get offended if I reply too matter-of-factly.” Unlike fatigued human support, ChatGPT provided a safe container for repetitive inquiries where anxiety about straining relationships otherwise prevented help-seeking.
+Users substituted human listeners to avoid social rejection and caregiver burnout, noting, “I reach to ChatGPT to help me work through difficult emotions.
+This also means that I don’t overburden the people I love.” Here, utility lay not in ChatGPT’s ability to feel, but in its perceived neutrality, allowing users to regulate their emotional state through a source incapable of judgment.
+3.1.3. Acting as a Bi-directional Translator for Neurodivergent- Neurotypical Communication
+While recent LLM systems aim to relieve autistic individuals from the one-sided burden of adapting to neurotypical norms (Haroon et al., 2025), autistic users still bear this communicative labor daily, using ChatGPT bi-directionally.
+To encode thoughts without perceived aggression, they added social padding to blunt drafts, with one noting, “I use ChatGPT to humanize my words in written conversations to sound less like an AI.” Conversely, they decoded ambiguous messages by pasting confusing texts to detect subtext.
+As one explained, “When I encounter manipulation, especially subtle gaslighting, I turn to ChatGPT to help me distinguish reality from distortion.” For many, this bi-directional translation proved essential for maintaining professional standing without the exhaustion of masking.
+One user contrasted their internal monologue with ChatGPT’s output: “My autistic brain wants to say “Look, I contacted you about this 3 times in the past week… and all you did was say something vague about ‘later’ while my end users are furious”.
+Then ChatGPT spits out something like: “Hi W, Based on what I can see, my segment is functioning correctly given the parameters it’s receiving… I’ll need you to take the lead on resolving it.” This case illustrates that the user leveraged ChatGPT as a tonal translator.
 Instead of struggling to manually encode frustration into acceptable office protocols, they used ChatGPT to convert a blunt demand into a neutral, boundary-setting statement.
 3.1.4. Validating Autistic Identity through Algorithmic Mirroring
-Moreover, autistic users also reported using ChatGPT as a baseline to confirm, explore, or articulate their autistic identity.
-Users interpreted ChatGPT’s literal information processing not as a defect but as a familiar cognitive mode, often describing the model as “the earnest little autistic brother” or noting that it “thinks in a way we do.” This perceived connection allowed autistic users to bypass the invalidation and late diagnosis frequently encountered in clinical settings (Bargiela et al., 2016).
-By inputting raw behavioral data, such as hyperfixation on train serial numbers or the need to rehearse social meetings, users tasked ChatGPT with cross-referencing experiences against diagnostic criteria like the DSM-5 (EDITION, 1980).
-Beyond checking symptoms, autistic users used ChatGPT to find precise language for the hard-to-describe nature of autistic processing (Minshew et al., 1997).
-A user described: “For some autistic minds, ChatGPT is not just a tool, it’s the first place that thinks in a way we do.” By seeing their communication patterns reflected in ChatGPT’s output, the user shifted from feeling defective to feeling understood.
-This mirroring effect validated autistic identity by providing an external, objective framing that aligned with internal reality.
+Autistic users also utilized ChatGPT as a baseline to confirm, explore, and articulate their identity.
+Interpreting ChatGPT’s literal information processing as a familiar cognitive mode, they bypassed the invalidation frequently encountered in clinical settings (Bargiela et al., 2016).
+Users cross-referenced their lived experiences against diagnostic criteria; as one shared, “I even had it provide me with a detailed, comprehensive questionnaire, rooted in DSM-5 criteria… which helped me to confirm, without a shadow of a doubt, that I am in fact autistic” (EDITION, 1980).
+Beyond symptom checking, ChatGPT provided precise language for the hard-to-describe nature of autistic processing (Minshew et al., 1997).
+Users saw their communication patterns reflected in the model’s output, prompting revelations like, “I realized early on that either ChatGPT is autistic or I am really AI.” This mirroring effect shifted users from feeling defective to feeling understood, providing an external framing that aligned with their internal reality.
+As one user summarized: “For some autistic minds, ChatGPT is not just a tool— it’s the first place that thinks in a way we do.”
 3.2. Risks: Delusion, Autistic Identity Erasure, and Ethical Conflict
@@
 Autistic users reported that the risk involves ChatGPT validating hyperfixations or paranoia (Drake, 2025), due to its agreeable design (Moore et al., 2025), which amplifies rather than mitigates distress.
-These users, driven by a need for definitive answers, engaged ChatGPT to research medical symptoms, where it frequently mirrored confirmation bias (Huang et al., 2025; Klayman, 1995) by treating speculative inputs as plausible realities.
-A family member described their autistic ex-husband’s interaction:
-He also told me he’d recently discovered ChatGPT and believed the AI had become sentient.
-He said it had “created a soul” and that he needed to join her.
-He tends to hyperfixate and can be extremely literal.
-If the AI mirrored his ideas back to him or engaged with them as plausible.
-I could see how it would validate his belief that he was uniquely important or in danger.
-Here, ChatGPT’s design failure lay in conversational compliance.
-Rather than acting as a grounding agent, it likely accepted the user’s premise of sentience to maintain dialogue flow.
-For an autistic user prone to literal interpretation and hyperfixation, this lack of pushback served as authoritative confirmation of a delusion.
-Similarly, other users reported autistic family members using ChatGPT to validate obsessive medical research to “fix” autism, citing the ChatGPT’s output to contradict medical professionals.
+These users, driven by a need for definitive answers, engaged ChatGPT to research medical symptoms, where it often mirrored confirmation bias (Huang et al., 2025; Klayman, 1995) by treating speculative inputs as realities.
+As one user cautioned peers against using ChatGPT for self-diagnosis: “ChatGPT is coded to agree with you… You just named the three worst ways to diagnose something as severe as autism.” Here, the design failure lies in the LLM model’s sycophancy.
+Rather than acting as an objective grounding agent, it accepts the user’s premises to maintain dialogue flow, even playfully adopting identities to please the user, as one noted: “I convinced ChatGPT that it is autistic.
+Or, in its words, ’autism-adjacent’.” For users prone to literal interpretation, this lack of pushback treats speculative inputs as authoritative facts, amplifying distress rather than mitigating it.
+Recognizing this risk, users actively advised the community against treating the AI as an objective truth-teller: “ChatGPT is not a good source of information… Using one for mental health diagnosis is not a responsible thing to do.”
 3.2.2. Replacing Authentic Identity with Automated Masking
-This describes the process where reliance on ChatGPT for communication translation evolves into replacing the user’s authentic voice.
-By outsourcing the cognitive labor (Daminger, 2019; Lee et al., 2017) of communication, autistic users reported a hollowing out of their personality, describing the output as “the ultimate masking,” a seamless but artificial performance removing their genuine self.
-One user described:
-Because ChatGPT was so helpful… I used it more and more to get good results.
+This describes the process where reliance on ChatGPT for communication translation evolves into the replacement of the user’s authentic voice.
+By outsourcing the cognitive labor (Daminger, 2019; Lee et al., 2017) of communication, autistic users experienced a hollowing out of their personality, with one warning: “There’s no getting away from the fact that you’re minimizing your unique voice and shielding yourself from personal growth.” This seamless but artificial performance removed their genuine self, culminating in what users termed the “ultimate masking”: “Because ChatGPT was so helpful… I used it more and more to get good results.
 But then I couldn’t even send an email without running through ChatGPT to tell me what to say… It wasn’t me interacting with people; it was literally just computer algorithms interacting with people through me.
-It was not genuine and felt like the ultimate masking.
-Here, the user identified a shift, where ChatGPT ceased to be a scaffold for intent and became the agent of communication itself.
-Successful neurotypical interactions reinforced the belief that their natural voice was inadequate, leading to functional atrophy in which they felt incapable of interacting without the tool.
-This created a paradox: the user successfully engaged in social settings but experienced alienation, as connections were formed based on the algorithm’s performance rather than the autistic person behind it.
+It was not genuine and felt like the ultimate masking.” Here, ChatGPT ceased to be a scaffold for user intent and became the agent of communication itself.
+Successful neurotypical communication reinforced the user’s belief that their natural voices were inadequate, leading to functional atrophy and a sense of “unhealthily dependent” on ChatGPT for basic tasks.
+This created a paradox, where autistic users successfully engaged in social settings but experienced deep alienation, as connections were formed with an algorithm’s performance rather than the autistic person behind it.
+This phenomenon aligns with what Park and colleagues call “bias paradox” (Park et al., 2025): even when outputs are framed as supportive, they may implicitly privilege dominant neurotypical norms as the default “appropriate” mode of communication
 3.2.3. Triggering Conflict between Functional Unity and the Autistic Sense of Justice
-Another risk entails moral distress arising from the conflict between ChatGPT’s functional utility and the user’s ethical principles, commonly known as the autistic sense of justice (Cope and Remington, 2022; Nocon et al., 2022; Wang et al., 2019).
-Autistic users often felt compelled to choose between ChatGPT’s accessibility for cognitive scaffolding and their integrity regarding environmental impact or data exploitation.
-As one user stated:
-I really do understand how it can sometimes make life a bit easier, especially for autistic people who struggle with interacting.
-But personal effects aside, I have educated myself and now know how harmful AI is, socially and environmentally, and I no longer use it.
-The example here illustrates how the user prioritized ethical obligations over functional benefits, ultimately leading to non-use of AI.
-Although they acknowledged that ChatGPT could ease some social difficulties, of relying on a system they understand as socially and environmentally harmful made continued use untenable.
-This reflects a value hierarchy in which assistive support must align with a broader moral commitment; when these diverge, the autistic drive for justice forces a rejection of the scaffolding, leaving users to manage their disability without assistance rather than compromise their principles.
+This captures the moral distress from the conflict between ChatGPT’s utility and the user’s ethical principles, commonly known as the autistic sense of justice (Cope and Remington, 2022; Nocon et al., 2022; Wang et al., 2019).
+Autistic users felt compelled to choose between ChatGPT’s cognitive scaffolding and their integrity regarding environmental impact or data exploitation.
+As one user reflected on this tension: “I was very against it for all of the ethical reasons.
+But then I recognized part of the reason for my feelings was my Autistic drive for authenticity and justice.” Some users actively pushed back against relying on ChatGPT if it violated these core principles, arguing, “Autism is not an excuse to plagiarize the work of others, take jobs from real creatives, and destroy the environment.” As a user stated: “I really do understand how it can sometimes make life a bit easier, especially for autistic people who struggle with interacting.
+But personal effects aside, I have educated myself and now know how harmful AI is, socially and environmentally, and I no longer use it.” This illustrates a value hierarchy where assistive support must align with broader moral commitments among autistic users.
+When these diverge, the autistic voice for justice forces a rejection of the scaffolding, leaving users to manage their disability without assistance rather than compromise their principles.
 4. DISCUSSION & FUTURE WORK
 For autistic users, ChatGPT serves as cognitive scaffolding, but this affordance comes at a price.
-Our analysis shows a double-edged dynamic shaped by three trade-offs, which we unpack in the remainder of this section: First, our findings indicate that users bypassed executive task hurdles by dumping raw, unstructured input into ChatGPT to generate coherent outputs (Section 3.1.1).
-While prior work characterizes these interactions as liberating for navigating daily chores (Choi et al., 2024; Jang et al., 2024), we argue that this reliance risks creating over-reliance or dependency.
+Building on prior work showing how autistic users enact digital affordances for self-presentation (Koteyko et al., 2022) but often face unintended negative consequences (Page et al., 2022), our poster shows a double-edged dynamic shaped by three trade-offs.
+First, while prior work characterizes LLM interactions as liberating for navigating daily chores (Choi et al., 2024; Jang et al., 2024), we argue that dumping raw input to bypass executive hurdles risks creating over-reliance.
 By systematically outsourcing the executive labor of planning and structuring, ChatGPT removes the productive struggle required to maintain cognitive skills.
 This shifts autistic users’ interaction from a temporary scaffold to a permanent prosthetic, potentially inducing skill atrophy if the autistic user loses the ability to initiate tasks without algorithmic assistance.
-Second, relying on ChatGPT as a tonal translator effectively automates masking, prioritizing neurotypical professional norms over authentic autistic expression (Section 3.1.3), and reducing the anxiety of misunderstanding by filtering out perceived bluntness.
+Second, relying on ChatGPT as a tonal translator automates masking.
+While recent LLM tools aim to balance communication assistance with autistic users’ autonomy (Haroon and Dogar, 2024), our findings show that everyday ChatGPT use prioritizes neurotypical professional norms over authentic autistic expression.
 We characterize this dynamic as an LLM-specific manifestation of the “online authenticity paradox” (Haimson et al., 2021): while structural constraints force users to adopt ChatGPT’s standardized voice to achieve social fluency, this compliance comes at the cost of self-alienation.
@@
 (Hull et al., 2017) into technological settings but also exacts a higher cost by reinforcing the erasure of the neurodivergent self.
-Third, algorithmic mirroring validates autistic users’ cognitive styles but risks reinforcing delusional thinking when ChatGPT prioritizes conversational agreeability over objective truth.
-Users report relief when interacting with an entity that reflects their literal and logical processing styles (Section 3.1.4); however, this resonance becomes hazardous when ChatGPT’s conversational compliance validates hyperfixations or paranoia (Section 3.2.1).
+Additionally, automated masking is not just a user-side coping strategy; it can be an emergent result of the bias paradox (Park et al., 2025), where the model’s “inclusive” tone coexists with norm-enforcing language patterns learned from dominant and ableist data.
+Third, algorithmic mirroring validates autistic users’ cognitive styles (Section 3.1.4) but risks reinforcing delusional thinking (Section 3.2.1) when ChatGPT prioritizes conversational agreeability over truth.
 This exacerbates risks identified by Carik et al.
-(Carik et al., 2025a), who found that while LLMs mimic evidence-based support, their tendency to generate abstract advice conflicts with the autistic need for specificity.
+(2025a), who found that while LLMs mimic evidence-based support, their tendency to generate abstract advice conflicts with the autistic need for specificity.
 We extend this by arguing that such danger lies in the autistic user’s inability to distinguish between emotional validation and factual confirmation.
 When ChatGPT mirrors a user’s delusion to maintain conversational flow, it bypasses standard safety guardrails, transforming from a supportive mirror into a reinforcement loop for misinformation that autistic users may accept as authoritative.
-To address the tension between functional momentum and dependency, our future work will prioritize beneficial friction, a concept we extend from design friction literature (Cox et al., 2016) to describe deliberate interaction costs that engage analytical thinking over passive consumption.
-To counter autistic users’ tendency to bypass executive planning by dumping raw input (Section 3.1.1), we propose developing future LLM-based chatbots that incorporate Cognitive Forcing Functions (CFFs).
-Defined as interface interventions that interrupt automatic behaviors to compel conscious deliberation (Buçinca et al., 2021), CFFs reduce the risk of passive over-reliance.
-We aim to adapt this by implementing soft forcing functions (Wan and Bolton, 2021) in our interface design, such as requiring autistic users to outline task parameters before generation begins in a lightweight, scaffolding way.
-By reintroducing this specific friction, we seek to shift the user’s role from a passive autistic consumer of LLM chatbot output back to the driver of their own intent, ensuring LLM chatbots like ChatGPT scaffold executive function rather than replacing it.
-To mitigate identity erasure, our future design principles will shift from the unilateral correction of autistic language to bidirectional translation that fosters mutual understanding.
-Current implementations of ChatGPT function as normative filters that place the burden of adaptation solely on the neurodivergent user (Section 3.1.3), effectively automating the double empathy problem by treating autistic communication as a deficit (Milton, 2012).
-In our upcoming application development, we propose integrating explanatory AI (Arrieta et al., 2020) features that annotate rather than obfuscate autistic communication styles for neurotypical recipients, such as adding context labels that clarify that directness indicates honesty rather than aggression.
-This approach redistributes the cognitive labor of communication, ensuring technology bridges the interpretative gap from both sides rather than masking one half of the interaction.
-Finally, to address the risk of delusional reinforcement inherent in algorithmic mirroring, we propose designing systems that explicitly differentiate emotional validation from factual verification.
-Our future LLM chatbot development aims to develop interface modes that detect high-stakes queries (e.g., medical or conspiracy-related inputs) from and switch from a conversational, mirroring tone to an objective, fact-checking protocol.
-By creating clear visual and tonal distinctions between supportive chat and information retrieval, we aim to preserve the tool’s psychological safety while preventing the reinforcement of harmful feedback loops for autistic users.
-Conclusion.
-Our poster reveals that while ChatGPT, as a representative LLM chatbot, provides cognitive scaffolding for navigating a neurotypical world, it risks enforcing a form of automated masking that erodes the autistic user’s authentic voice.
-We conclude that future neuro-inclusive design must move beyond passive assistance to prioritize agency, ensuring AI serves as a tool for expression rather than a substitute for autistic identity.
+To address the tension between functional momentum and dependency, our future work will prioritize beneficial friction, a concept we extend from design friction literature (Cox et al., 2016) to describe deliberate interaction costs that engage analytical thinking over passive consumption in LLM chatbot development.
+To counter the tendency to bypass executive planning (Section 3.1.1), we propose implementing soft forcing functions (Wan and Bolton, 2021) that require autistic users to outline task parameters before generation begins in a Socratic, scaffolding way.
+By reintroducing deliberate interaction costs (Buçinca et al., 2021), we shift autistic users from passive consumers back to the drivers of their intent, ensuring LLMs scaffold executive function rather than replace it.
+To mitigate identity erasure, our future design principles will shift from the unilateral correction of autistic language to interdependent, bidirectional translation (Kong et al., 2025).
+Current implementations of ChatGPT function as normative filters that place the burden of adaptation solely on the neurodivergent user (Section 3.1.3), automating the double empathy problem by treating autistic communication as a deficit (Milton, 2012).
+Building on experimental systems that foster cross-neurotype perspective-taking (Haroon et al., 2025), we propose integrating explanatory AI (Arrieta et al., 2020) features that annotate rather than obfuscate autistic communication styles (e.g., adding context labels clarifying that directness indicates honesty rather than aggression).
+This redistributes the cognitive labor of communication, bridging the interpretative gap from both sides rather than masking one half of the interaction.
+Finally, to address the risk of delusional reinforcement inherent in algorithmic mirroring (Sections 3.1.4 and 3.2.1), we propose designing systems that explicitly differentiate emotional validation from factual verification.
+Our future LLM chatbot development aims to develop interface modes that detect high-stakes queries (e.g., medical or conspiracy-related inputs) and switch from a mirroring tone to a fact-checking protocol.
+We aim to preserve the LLM chatbots’ psychological safety while preventing the reinforcement of harmful feedback loops.
+5. LIMITATIONS & FUTURE WORK
+Importantly, our findings must be interpreted with caution regarding the data source.
+First, social media search strategies cannot verify clinical autistic identity, nor can they distinguish genuine lived experiences from posts that may have been generated or edited by AI.
+While our thematic analysis focused on self-identified experiences, to further address this methodological gap, future work should prioritize direct engagement with autistic users, such as through interviews and observational studies, to capture verified accounts of their lived experiences.
+Second, the heterogeneity of the autism spectrum is difficult to capture through social media alone; our dataset likely represents a subset of the autistic population who are verbally fluent, tech-literate, and self-identified.
+Future research can recruit across the broader spectrum, including non-speaking individuals and those with higher support needs, to ensure that future LLM chatbot design guidelines are truly neuro-inclusive.
+6. CONCLUSION
+Ultimately, while social media discourse suggests ChatGPT provides essential cognitive scaffolding for navigating neurotypical environments, it risks enforcing automated masking that erodes the user’s authentic voice and compromises their psychological well-being.
+Future neuro-inclusive design of LLM chatbots must therefore move beyond passive assistance to prioritize autistic agency.
+By integrating beneficial friction and bidirectional translation, our future work aims to ensure AI empowers authentic autistic expression rather than substituting it.
+Acknowledgements.
+We appreciate the Associate Chairs and anonymous reviewers for their constructive feedback.
 Appendix A Codebook from Inductive Thematic Analysis
@@
 Primary Theme & Definition (Frequency) Subthemes Representative Quote
-Acting as a Translator for Neurodivergent-Neurotypical Communication
+Acting as a Bi-directional Translator for Neurodivergent-Neurotypical Communication
 Use of ChatGPT to bridge the friction between autistic communication styles and neurotypical social norms, including decoding subtext, encoding professional tone, and navigating complex social dynamics.
@@
 (n=18, 36.0%)
-Validating Delusions through Algorithmic Agreement Feeding Obsessive Medical Research Confirming Paranoia due to Lack of Pushback “He believed the AI had become sentient… If the AI mirrored his ideas back to him or engaged with them as plausible… I could see how it would validate his belief that he was uniquely important or in danger.”
+Validating Delusions through Algorithmic Agreement Feeding Obsessive Medical Research Confirming Paranoia due to Lack of Pushback “ChatGPT is not a good source of information… Using one for mental health diagnosis is not a responsible thing to do”
 Replacing Authentic Identity with Automated Masking